Louise is painting around a window frame on the outside of her house. Her ladder is 3.5m long and she placed it 1 m from the house. What is the vertical distance that the ladder reaches up the house? Round to one decimal place. Use the Pythagorean theorem.
1
Expert's answer
2020-05-04T18:39:06-0400

Consider the ladder leans against the wall of the house as shown in the figure below:





The figure above shown is a right angled triangle.


Here, "a=1,b=y,c=3.5"


Recall the Pythagoras theorem "c^2=a^2+b^2" .


Plug "a=1,b=y,c=3.5" into above formula as,


"3.5^2=1^2+y^2"

"12.25=1+y^2"


Subtract "1" from both sides to isolate the variable "y" as,


"12.25-1=1+y^2-1"

"11.25=y^2"

"y^2=\frac{45}{4}"


Rewrite the fraction "\frac{45}{4}" into perfect square as,


"y^2= (\frac{3\sqrt {5}}{2})^2"


Using property of radicals, if "x^2=a^2" , then "x=\pm \sqrt {a^2}=\pm a" , therefore, the value of "y" is,


"y=\pm \frac{3\sqrt{5}}{2}"


Since, negative value has no significant as it is a distance, so negative value neglected.


So, the value of "y" is "y=\frac{3\sqrt{5}}{2}"


Therefore, the vertical distance that the ladder reaches up the house is "y=\frac{3\sqrt{5}}{2} \approx 3.4" m.